OpenStudy (anonymous):

What should be added to both sides of this equation to solve for the variable?r + (-12) = 19

OpenStudy (anonymous):

12 should be added to both sides

OpenStudy (anonymous):

this one is ganna be postive 12 right

OpenStudy (anonymous):

\[r+-12=12\] \[r+-12+12=12+12\] r=24

OpenStudy (anonymous):

i got it

OpenStudy (anonymous):

your welcome medal if i did well :D